How do I plug in a translated hyperbola equation into my graphing calculator? For example, (y-3)^2/(4) - (x+1)^2/(9) = 1 ?
Answer by ewatrrr(24785)

 
Hi,
(y-3)^2/(4) - (x+1)^2/(9) = 1
y = 3 +(2)sqrt(1+(x+1)^2/9)
y = 3 -(2)sqrt(1+(x+1)^2/9)